OpenStudy (anonymous):

Please help. How do you simplify (sint/(1+cost))+((1+cost)/sint) to 2csct?

OpenStudy (anonymous):

\[\frac{ \sin t }{ 1+\cos t }+\frac{ 1+\cos t }{ \sin t }=2\csc t\]

OpenStudy (anonymous):

On the left. Find a common denominator then add the top

OpenStudy (anonymous):

I got to (sint/(1+cost))+((1+cost)/sint)

OpenStudy (cggurumanjunath):